OpenStudy (anonymous):

How to do you graph 3c + 2d = -6 and -3c + 2d = 6 as a slope intercept? Please help!!

OpenStudy (anonymous):

in slope intercept form d=-3 -(3/2)c

OpenStudy (anonymous):

3c + 2d= -6 -3c + 2d =6 change both equation to slope intrcept form. . 2d=-3c-6 2d= 3c+6 d=-3/2c-3 d=3/2c+3 therfor the slope are -3/2 and 3/2 since we have the value of the y intercept - -3 and 3 threfor, we can create points (0,-3) and (0,3) to graph use the follwing: m=-3/2 (0,-3) m=3/2 (0,3) I THINK THEY ARE PARALLEL. . .